6. The mean waiting time at the drive-through of a fast-food restaurant from the time an order is placed to the time the order is received is 84.3 seconds. A manager devises a new drive-through system that he believes will decrease the wait time. He initiates the new system at his restaurant and measures the wait time for 10 randomly selected orders. The wait times are in the following table: 108.5 67.4 58.0 75.9 65.1 80.4 95.5 86.3 70.9 72.0 a) Find the sample mean and standard deviation using Minitab. b) Is this data normally distributed? Create a normal probability plot on Minitab. c) is this new system effective? Conduct the approprlate test at the a 0.1 level of significance.
Answer #1

Ans:

a)

n=10

sample mean=78

standard deviation=15.205

b)

Nornal Probability Plot 1.64 1.14 0.64 0.14 N 52.95 57.95 62.95 67.95 72.95 77.95 82.95 87.95 92.95 97.95 102.95 107.95 112.95 0.36 0.86 1.36

c)

Test statistic:

t=(78-84.3)/(15.205/sqrt(10))

t=-1.310

df=10-1=9

p-value=tdist(1.310,9,1)=0.1113

As,p-value>0.1,we fail to reject the null hypothesis.

There is not sufficient evidence to conclude that waiting time is decreased(i.e. new system is effective)